# Not All Diffusion Model Activations Have Been Evaluated as Discriminative Features

### v1

@@ 2 Related Work
The diffusion feature approach has seen various improvements. Some techniques toggle the input settings of diffusion models, such as seeking better timesteps [57, 28]. The others add trainable parameters outside the diffusion model to refine the outputs or provide an efficient fine-tuning alternative [55, 26, 43]. Additionally, some studies explore completely training-free methods that utilize spatial attention information [48, 40].
@@ 3 Preliminaries: Architecture of Diffusion U-Nets
Furthermore, each ViT module consists of several stacked basic blocks. A basic block typically has a self-attention layer to perform attention on the image itself and a feed-forward layer, essentially a two-layer MLP [14]. The diffusion U-Net introduces an additional cross-attention layer between the two layers, enabling the fusion of the image and additional textual prompts. Besides, each layer includes a residual connection, meaning that the increment activation added to the outer residual is also the internal residual.
@@ 4 Distinct Properties of Diffusion U-Nets
The diffusion U-Net has many interesting properties, but now we only focus on those distinct from the knowledge of traditional U-Net [35] or ViTs [41]. As shown in Figure 3, this section highlights three noticeable properties, each of which corresponds to a different level of the diffusion U-Net architecture described in Section 3. Notably, these properties can be universally observed in different samples and diffusion models, though all visualization in the main content is conducted on the same image and SDXL for consistency. Additional visualization is provided in Appendix A. Besides, the omitted properties are available in Appendix E.
@@ 4.1 Asymmetric Diffusion Noises
How does the influence of diffusion noises spread across diffusion U-Nets? As shown in Figure 3(a), diffusion noises exist throughout the entire down-stage, with a decreasing magnitude. Remarkably, during the early half of the up-stage, activations are rather clean, with no perceivable diffusion noises. Only in the latter half do diffusion noises start to resurface. As shown in Appendix A, such an asymmetric pattern is consistent across activations in different diffusion models. It is noticeable that, although being universal among diffusion U-Nets, this property was not reported in the research of traditional U-Nets [54], to the best of our knowledge.
@@ 4.2 In-Resolution Granularity Changes
The second property is at the in-resolution level and closely related to recent advances in diffusion architectures. Specifically, the design of U-Net follows the idea of resolution hierarchy [35]. Consequently, the overall architecture displays a fine-coarse-fine granularity trend, looking like the alphabet “U”. Traditional U-Nets implement this architecture with a relatively large number of resolutions, while each resolution is typically small, equipped with two or three simple convolutional layers. Hence, the research of traditional U-Net focuses on the granularity changes across resolutions, implicitly assuming that the change within a single resolution is negligible.
@@ 4.3 Locality without Positional Embeddings
In contrast, ViT modules in diffusion U-Nets do not use positional embeddings [34, 33], perhaps because the other U-Net components have provided sufficient spatial cues. This change results in distinct properties of the activations. On one hand, positional information is negligible for most activations despite how near they are to the inputs. On the other hand, the queries and keys of self-attention still display non-negligible positional information, marked with orange circles in Figure 3(c). Specifically, the latent pixel on the horse’s neck is a light blue color, similar to the pixels to its left that actually represent the background. In contrast, the pixels above the circle are in purple color, though they also represent the horse’s neck. Such comparison shows that a latent pixel is more similar to other pixels that are spatially near it than those semantically closer to it.
We name this phenomenon locality since it has a different mechanism from that induced by positional embeddings. As pointed out in [11], self-attention allows ViT to integrate global and local information even in the shallow layers, and the attention scope enlarges w.r.t. layer depth. Even without positional embeddings, self-attention activations are generally consistent with this insight, but the magnitude has greatly reduced. As shown in Figure 3(c), in shallow activations, locality exists but is inherently weaker than that induced by positional embeddings, such that more semantic information is reserved. As the layer goes deeper, locality quickly diminishes.
@@ 4.4 Universality of Three Properties
Although the visualization in Figure 3 is conducted only on SDXL, the scope of these properties is not limited to the specific architecture. Further evidence supporting the following claims is available in Appendix A.
@@
    In-resolution granularity changes come from the “fatness” of U-Nets, making it applicable to more traditional U-Nets.
@@ 5.2 Quantitative Comparison
Given the capability ranking, feature selection is in fact flexible, as it is possible to combine multiple activations and specifically tune the choice for a task. For practicality, we provide generic combinations of features for SDv1.5 and SDXL in Appendix B which are likely to perform well generally, according to the results detailed in Appendix D. Each of them consists of four activations mainly from ViT modules rather than inter-module positions. Taking SDv1.5 as an example, one of the four selected features is from one self-attention layer in the highest resolution, which utilizes our observation in Section 4.
@@ 6.2 Empirical Results on Semantic Segmentation
Unlike the results in semantic correspondence, Legacy-XL outperforms Legacy-v1.5 on CityScapes (71.67 v.s. 64.01), and the performance gap between Legacy-v1.5 and Ours-v1.5 is narrow (64.01 v.s. 64.10). This is because this task utilizes a large-scale downstream model, which can significantly refine the input features and thus reduce the gap in feature quality. Nevertheless, Ours-XL still achieves a significant improvement over Legacy-XL (74.47 v.s. 71.67).
@@ 7 Conclusion and Future Work
However, we are not sure whether our observations can generalize well to recently-developed DiT models [32] since it has a markedly different architecture from diffusion U-Nets. Thus, analyzing DiT models is a promising topic for future research.
@@ A.2 Visualization from Other Diffusion Models
Conventional U-Net Activations. As comparison to diffusion U-Net, we also visualize some activations from a conventional U-Net in fig. 15, where the three properties of diffusion U-Net can hardly be observed.
@@ Appendix B Details of Our Method
In this section, we index all diffusion U-Net components in the order in which they are activated during a network forward run. Besides descriptions in natural language, we also index the activations using the same notations as used in our code implementation.
@@ Appendix F Future Direction
It might be a good direction to next focus on more challenging discrimination scenarios. Specifically, long tail and out-of-distribution are important problems in discrimination, which can greatly hinder the performance of models that work well under i.i.d. settings [46]. There have been a few attempts on addressing this more challenging problem with diffusion models. For example, the disentanglement property of prompts might grant diffusion features cross-domain capability [15]. It is also possible to utilize diffusion models to synthesize training samples to adjust training data distribution [47, 52]. Given these attempts, more efforts are still required to be put into this direction. Moreover, there is a recent study that might boost long tail and out-of-distribution discrimination using diffusion models [16]. To be specific, AUC is an evaluation metric as well as loss function that promotes good performance on both head and tail samples, which is an important tool for long tail and out-of-distribution problems [44, 45]. Conventionally, AUC is applicable to image classification but not semantic segmentation and other pixel-level tasks, but the aforementioned study manages to adopt AUC for semantic segmentation. Given this new tool, it is now made more viable to attempt to enhance diffusion feature on long tail and out-of-distribution problems.

### v2

@@ 2 Related Work
The diffusion feature approach has seen various improvements. Some techniques toggle the input settings of diffusion models, such as seeking better timesteps [60, 29]. The others add trainable parameters outside the diffusion model to refine the outputs or provide an efficient fine-tuning alternative [58, 27, 46]. Additionally, some studies explore completely training-free methods that utilize spatial attention information [51, 43], and some attempts focus on text-free diffusion models [32, 33].
@@ 3 Preliminaries: Architecture of Diffusion U-Nets
Furthermore, each ViT module consists of several stacked basic blocks. A basic block typically has a self-attention layer to perform attention on the image itself and a feed-forward layer, essentially a two-layer MLP [15]. Modern diffusion U-Net introduces an additional cross-attention layer between the two layers, enabling the fusion of the image and additional textual prompts. Besides, each layer includes a residual connection, meaning that the increment activation added to the outer residual is also the internal residual.
@@ 4 Distinct Properties of Diffusion U-Nets
The diffusion U-Net has many interesting properties, but now we only focus on those distinct from the knowledge of traditional U-Net [38] or ViTs [44]. As shown in Figure 3, this section highlights three noticeable properties, each of which corresponds to a different level of the diffusion U-Net architecture described in Section 3. Notably, these properties can be universally observed in different samples and diffusion models, though all visualization in the main content is conducted on the same image and SDXL model for consistency. Additional visualization is provided in Appendix A. Besides, the omitted properties are available in Appendix E.
@@ 4.1 Asymmetric Diffusion Noises
How does the influence of diffusion noises spread across diffusion U-Nets? As shown in Figure 3(a), diffusion noises exist throughout the entire down-stage, with a decreasing magnitude. Remarkably, during the early half of the up-stage, activations are rather clean, with no perceivable diffusion noises. Only in the later half do diffusion noises start to resurface. The existence of this asymmetric behavior can also be indirectly supported by the ablation curves in many fellow studies such as [2, 32, 33]. Furthermore, as shown in Appendix A, such an asymmetric pattern is consistent across activations in different diffusion models.
@@ 4.2 In-Resolution Granularity Changes
The second property is at the in-resolution level and closely related to recent advances in diffusion architectures. Specifically, the design of U-Net follows the idea of resolution hierarchy [38]. Consequently, the overall architecture displays a fine-coarse-fine granularity trend, looking like the alphabet “U”. Traditional U-Nets implement this architecture with a relatively large number of resolutions, while each resolution is typically small, equipped with two or three simple convolutional layers. Hence, the understanding of traditional U-Net focuses on the granularity changes across resolutions, implicitly assuming that the change within a single resolution is negligible.
@@ 4.3 Locality without Positional Embeddings
In contrast, ViT modules in diffusion U-Nets do not use positional embeddings [37, 36], perhaps because the other U-Net components have provided sufficient spatial cues. This change results in distinct properties of the activations. On one hand, positional information is negligible for most activations despite how near they are to the inputs. For example, even in the first basic block, cross-attention query activations contain no perceivable positional information. On the other hand, the queries and keys of self-attention still display non-negligible positional information, marked with orange circles in Figure 3(c). Specifically, the latent pixel on the horse’s neck is a light blue color, similar to the pixels to its left that actually represent the background. In contrast, the pixels above the circle are in purple color, though they also represent the horse’s neck. Such comparison shows that a latent pixel is more similar to other pixels that are spatially near it than those semantically closer to it.
We name this phenomenon locality since it has a different mechanism from that induced by positional embeddings. As pointed out in [12], self-attention allows ViT to integrate global and local information even in the shallow layers, and the attention scope enlarges w.r.t. layer depth. Even without positional embeddings, self-attention activations are generally consistent with this insight, leading to the existence of locality. Nevertheless, the magnitude has indeed greatly reduced, compared to the visualization of conventional ViT activations in  [1]. As shown in Figure 3(c), in shallow activations, locality exists but is inherently weaker, as much semantic information is still reserved. In addition, locality quickly diminishes as the layer goes deeper.
@@ 4.4 Universality of Three Properties
Although the visualization in Figure 3 is conducted only on SDXL, the scope of these properties is not limited to the specific architecture. Further supporting evidence is available in Appendix A.
@@
    In-resolution granularity changes come from the “fatness” of U-Nets, making it potentially applicable to more traditional U-Nets.
@@ 5.2 Quantitative Comparison
Given the capability ranking, feature selection is in fact flexible, as it is possible to combine multiple activations and specifically tune the choice for a task. For practicality, we provide off-the-shelf feature combinations for SDv1.5 and SDXL that are likely to generically perform well in Appendix B, according to the results detailed in Appendix D. Each of them consists of four activations mainly from ViT modules rather than inter-module positions. Taking SDv1.5 as an example, one of the four selected features is from one self-attention layer in the highest resolution, which utilizes our observation in Section 4.
@@ 6.2 Empirical Results on Semantic Segmentation
Unlike the results in semantic correspondence, Legacy-XL outperforms Legacy-v1.5 on CityScapes (71.67 v.s. 64.01), and the performance gap between Legacy-v1.5 and Ours-v1.5 is narrow (64.01 v.s. 64.10). This is because this task utilizes a relatively large-scale downstream model, which can significantly refine the input features and thus reduce the gap in feature quality. Nevertheless, Ours-XL still achieves a significant improvement over Legacy-XL (74.47 v.s. 71.67).
@@ 7 Conclusion and Future Work
However, we are not sure whether our observations can generalize well to recently-developed DiT models [35] since they have a markedly different architecture from U-Net-based diffusion models. Thus, analyzing DiT models is a promising topic for future research.
@@ A.2 Visualization from Other Diffusion Models
Conventional U-Net Activations. As comparison to diffusion U-Net, we also visualize some activations from a conventional U-Net in Figure 15, where the three properties of diffusion U-Net can hardly be observed.
@@ Appendix B Details of Our Method
In this section, we index all diffusion U-Net components in the order in which they are activated during a network forward run. Besides descriptions in natural language, we also index the activations using the same notations as used in our code implementation. These notations denote stages using down/mid/up, resolutions as level, and module index as repeat. Index starts to count from 0 instead of 1, following the convention in coding.
@@ Appendix F Future Direction
It might be a good future direction to focus on more challenging discrimination scenarios. Specifically, long tail and out-of-distribution are important problems in discrimination, which can greatly hinder the performance of models that work well under i.i.d. settings [49]. There have been a few attempts to address this more challenging problem with diffusion models. For example, the disentanglement property of prompts might grant diffusion features cross-domain capability [16]. It is also possible to utilize diffusion models to synthesize training samples to adjust training data distribution [50, 55]. Given these attempts, more efforts are still required to be put in this direction. Moreover, there is a recent study [17] that might boost long tail and out-of-distribution discrimination using diffusion models. To be specific, AUC is an evaluation metric as well as a loss function that promotes good performance on both head and tail samples, which is an important tool for long tail and out-of-distribution studies [47, 48]. Conventionally, AUC is applicable to image classification but not semantic segmentation and other pixel-level tasks, but the aforementioned study manages to adopt AUC for semantic segmentation. Given this new tool, it is now made more viable to attempt to enhance diffusion feature on long tail and out-of-distribution problems.